OpenStudy (anonymous):

25*2-7squared/(2squared-(-7))squared

OpenStudy (anonymous):

No, the whole bottom part is squared, not just the seven.

OpenStudy (mertsj):

\[\frac{25(2)-7^{2}}{[2^{2}-(-7)]^{2}}\]

OpenStudy (mertsj):

\[\frac{1}{[4+7]^{2}}\]

OpenStudy (mertsj):

\[\frac{1}{121}\]